# Bogomolov property and Galois representations

### v1

@@ 2. $\bullet$
(see the proof of Lemma 4.7), its image coincides with $N_{\mathbb{Q}_{q}/\mathbb{Q}_{p}}(1+p^{n-1}\mathbb{Z}_{q})$ mod $1+p^{n}\mathbb{Z}_{p}$, and the norm map from $1+p^{n-1}\mathbb{Z}_{q}\rightarrow 1+p^{n-1}\mathbb{Z}_{p}$ is surjective (see for example [24, Ch. V, §2, Prop. 3]). Therefore also the trace from $\mathcal{L}_{n}(\overline{D_{n}})$ to ${\mathbb{F}}_{p}$ is surjective.
Thus $\mathcal{L}_{n}(\overline{D_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$ as claimed.
6. Proof of Theorem 1.3
@@
We want to apply Proposition 2.6 with $L={\mathbb{Q}}(p^{\infty})$. We first check Assumption 2.2. By Lemma 4.7 (iii) (with $M=2$) there exists an element $\tau\in Gal(\mathbb{Q}_{q}(p^{\infty})/\mathbb{Q}_{q})$ whose image by $\rho_{v}$ is the scalar matrix $\left(\begin{smallmatrix}2&0\\ 0&2\end{smallmatrix}\right)$. Since $\rho_{v}|_{G_{\mathbb{Q}_{q}}}$ is totally ramified and $\mathbb{Q}_{q}/\mathbb{Q}_{p}$ is unramified, $\tau$ can be lifted to an element in the inertia subgroup $I_{p}$ of $G_{{\mathbb{Q}}_{p}}$. Indeed $\mathbb{Q}_{q}(p^{\infty})$ and $\mathbb{Q}_{q}^{nr}$ are linearly dijoint over $\mathbb{Q}_{q}$ and $I_{p}=Gal(\overline{\mathbb{Q}}_{p}/\mathbb{Q}_{q}^{nr})$ (where $\mathbb{Q}_{q}^{nr}$ denotes the maximal unramified extension of $\mathbb{Q}_{p}$), therefore $\tau$ can be lift to an element in $I_{p}$. As its image is a scalar, $\tau$ lies in the center of $\mathrm{Gal}(L/{\mathbb{Q}})$. Its action on the $p^{\infty}$-th roots of unity is given by the $p$-adic cyclotomic character, i.e. the determinant of $\rho_{v}(\tau)$, which is $4$. Then $\tau(\zeta)=\zeta^{4}$ for every $\zeta\in{\boldsymbol{\mu}}_{p^{\infty}}$.
@@
We now check Assumption 2.3. We consider the filtration given by $L_{n}={\mathbb{Q}}(p^{n})$. Let $n\geq 1$. Fix a place of ${\mathbb{Q}}(p^{\infty})$ over $p$ corresponding to an embedding ¯¯$\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}}_{p}$. Then, by Lemma 4.7 (ii) and in the notation of Assumption (2.3), $e_{n}=(q-1)q^{n-1}$, $i_{n}=q^{n-1}-1$ and $G^{n}_{i_{n}}\simeq\mathrm{Gal}(\mathbb{Q}_{q}(p^{n})/\mathbb{Q}_{q}(p^{n-1}))$. Thus $e_{n}/(i_{n}+1)=q-1\leq p^{2}$. As already remarked $\rho(p^{n})$ identifies ${\overline{D_{n}}}=H_{n}$ to a subgroup of
$1+p^{n-1}\mathrm{M}_{2}(\mathcal{O}/(p^{n}))\simeq\mathrm{M}_{2}(\mathcal{O}/(p))$
so that $|H_{n}|\leq p^{4[K:{\mathbb{Q}}]}$, where we recall that $K$ is the number field generated by the coefficients of our modular form. Moreover, by Proposition 5.1, $L_{n}^{H_{n}}=L_{n-1}$. The Assumption 2.3 therefore holds with $C_{1}=p^{2}$ and $C_{2}=p^{4d}$, where $d=[K:\mathbb{Q}]$.
@@ Appendix A On the linear groups of a finite product of local rings
Let ${\mathcal{A}}=\prod_{i=1}^{s}{\mathcal{A}}_{i}$ be a finite product of local $\mathbb{F}_{p}$-algebras. In this appendix we bring together some results on the linear groups overs ${\mathcal{A}}$.
@@ A.1. A result on the normal closure
We define
$\widehat{G}({\mathcal{A}})=\{\gamma\in\mathrm{GL}_{2}({\mathcal{A}})\ |\ \det(\gamma)\in\mathbb{F}_{p}^{\times}\}$
It is the semidirect product $\mathrm{SL}_{2}({\mathcal{A}})\rtimes\mathbb{F}_{p}^{\times}$, where $x\in\mathbb{F}^{\times}$ acts on $\mathrm{SL}_{2}({\mathcal{A}})$ by conjugation by $\left(\begin{smallmatrix}x&0\\ 0&1\end{smallmatrix}\right)$. Note that $\widehat{G}(\mathbb{F}_{p})=\mathrm{GL}_{2}(\mathbb{F}_{p})$. The aim of this subsection is to prove the following proposition.
@@
Let us assume $p\geq 5$. Let $H$ be a subgroup of $\mathrm{GL}_{2}(\mathbb{F}_{p})$ such that $|H|>2(p-1)$ and such that the determinant maps surjectively onto $\mathbb{F}_{p}^{\times}$. Then the normal closure of $H$ in $\widehat{G}({\mathcal{A}})$ is $\widehat{G}({\mathcal{A}})$ itself.
@@
Let ${\mathcal{B}}$ be the $\mathbb{F}_{p}$-vector subspace of ${\mathcal{A}}$ generated by the squares of all the invertible elements of ${\mathcal{A}}$. Note that ${\mathcal{B}}$ is a subring with unity of ${\mathcal{A}}$ which contains $\mathbb{F}_{p}$. For every $i=1,\ldots,k$ denote by $e_{i}\in{\mathcal{A}}$ the idempotent projecting to the summand ${\mathcal{A}}_{i}$. Firstly we show that $e_{i}\in{\mathcal{B}}$ for every $i$. Indeed, choose two elements $a$, $b$ in $\mathbb{F}_{p}^{\times}$ such that $a^{2}\neq b^{2}$ (which is possible because $|\mathbb{F}_{p}^{\star}|>2$) and let $\alpha=ae_{i}+\sum_{j\not=i}e_{j}$, $\beta=be_{i}+\sum_{j\not=i}e_{j}$. Then $\alpha^{2}-\beta^{2}=(a^{2}-b^{2})e_{i}$ is a non-zero $\mathbb{F}_{p}$-multiple of $e_{i}$ belonging to ${\mathcal{B}}$.
@@
Let us assume $p\geq 5$. Let $H$ be a subgroup of $\mathrm{GL}_{2}(\mathbb{F}_{p})$ such that $|H|>2(p-1)$ and such that the determinant maps surjectively onto $\mathbb{F}_{p}^{\times}$. Then the normal closure of $H$ in $\mathrm{GL}_{2}(\mathbb{F}_{p})$ is $\mathrm{GL}_{2}(\mathbb{F}_{p})$ itself.
@@
Let $N$ be the normal closure of $H$ in $\mathrm{GL}_{2}(\mathbb{F}_{p})$. Since $|N|>2(p-1)$, the intersection $N\cap\mathrm{SL}_{2}(\mathbb{F}_{p})$ is a non trivial normal subgroup of $\mathrm{SL}_{2}(\mathbb{F}_{p})$ of order strictly greater than $2$. Since $|\mathbb{F}|\geq 4$, the only proper normal subgroup of $SL_{2}(\mathbb{F})$ is {±$\{\pm Id\}$ (see for example [26, Theorem 8.3 and Lemma 8.6, p.539]). Thus $\mathrm{SL}_{2}(\mathbb{F}_{p})\subseteq N$. Since the projection over $\mathbb{F}_{p}^{\times}$ is surjective, the result follows. ∎
We can now prove Proposition A.1.1. Let $N$ be the normal closure of $H$ in $\widehat{G}({\mathcal{A}})$. By Lemma A.1.4, $N$ contains $\mathrm{GL}_{2}(\mathbb{F}_{p})$, so that it contains $\mathrm{SL}_{2}({\mathcal{A}})$ by Lemma A.1.3. The surjectivity of the determinant then shows that $N=\widehat{G}({\mathcal{A}})$.
@@ A.2. Irreducibility of the adjoint representation
By (A.1), any invariant subspace containing one of $A,B,C$ will also contain the others. Given some arbitrary (non-zero) matrix $M=aA+bB+cC$ of trace 0, put
@@
We can now prove Proposition A.2.5. Firstly we notice that $\mathrm{M}_{2}(\mathbb{F}_{p})^{0}\subseteq V$ by Lemma A.2.6. Since $M_{2}({\mathcal{A}})^{0}=M_{2}(\mathbb{F}_{p})^{0}\otimes_{\mathbb{F}}{\mathcal{A}}$, it suffices to show that $V$ is closed under the multiplication by elements in ${\mathcal{A}}$. Let $\alpha$ be an invertible element of ${\mathcal{A}}$ and let $B=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$. We have

### v3

@@ 2. Notations and preliminary results
(see the proof of Lemma 4.7), its image coincides with $N_{\mathbb{Q}_{q}/\mathbb{Q}_{p}}(1+p^{n-1}\mathbb{Z}_{q})$ mod $1+p^{n}\mathbb{Z}_{p}$, and the norm map from $1+p^{n-1}\mathbb{Z}_{q}\rightarrow 1+p^{n-1}\mathbb{Z}_{p}$ is surjective (see for example [26, Ch. V, §2, Prop. 3]). By condition 1.3 $(P3)$ $p\nmid k-1$ and thus the $(k-1)$-th power of the cyclotomic character is surjective. Therefore also the trace from $\mathcal{L}_{n}(\overline{H_{n}})$ to ${\mathbb{F}}_{p}$ is surjective.
Thus $\mathcal{L}_{n}(\overline{H_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$ as claimed.
Remark 5.2.
We have seen in the previous proof that $\overline{H_{n}}=\Omega_{n}$ can be identified to a subgroup of $\mathrm{GL}_{2}(\mathcal{O}/p\mathcal{O})$ if $n=1$ and to a subgroup of $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})$ if $n\geq 2$. Thus $|{\overline{H_{n}}}|\leq p^{4[{\mathbb{K}}:{\mathbb{Q}}]}$, where we recall that $\mathbb{K}$ is the number field generated by the coefficients of our modular form.
6. Proof of the main results.
$\rho_{p}=\prod_{v|p}\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\widehat{\mathcal{O}})$
and
$\rho^{(p)}=\prod_{v\nmid p}\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\widehat{\mathcal{O}}).$
Proposition 6.1.
If Assumption 1.3 $(P0)$ and $(P1)$ hold, then $\rho^{(p)}$ has property (B).
Proof.
Let ${\mathbb{F}}=\mathbb{Q}(\rho^{(p)})$. By our assumptions and by (2.1) and (2.2), ${\mathbb{F}}$ is unramified at $p$ and the characteristic polynomial of $\rho^{(p)}(\mathrm{Frob}_{p})$ is $X^{2}+p^{k-1}$. It follows that $\mathrm{Frob}_{p}^{2}$ lies in the center of $\mathrm{Gal}({\mathbb{F}}/{\mathbb{Q}})$. Then Property (B) for ${\mathbb{F}}$ is ensured by Proposition 3.1. ∎
Proof of Theorem 1.4.
@@
We want to apply Proposition 3.4 with ${\mathbb{L}}={\mathbb{Q}}(p^{\infty})$. We first check Assumption 3.2. By Lemma 4.7 (iii) (with $M=2$) there exists an element $\tau\in\mathrm{Gal}(\mathbb{Q}_{q}(p^{\infty})/\mathbb{Q}_{q})$ whose image by $\rho_{v}$ is the scalar matrix $\left(\begin{smallmatrix}2^{k-1}&0\\ 0&2^{k-1}\end{smallmatrix}\right)$. Since $\rho_{v}|_{G_{\mathbb{Q}_{q}}}$ is totally ramified and $\mathbb{Q}_{q}/\mathbb{Q}_{p}$ is unramified, $\tau$ can be lifted to an element in the inertia subgroup of $G_{\mathbb{Q}_{p}}$. As its image is a scalar, $\tau$ lies in the center of $\mathrm{Gal}({\mathbb{L}}/{\mathbb{Q}})$. Its action on the $p^{\infty}$-th roots of unity is given by the $p$-adic cyclotomic character, i.e. the determinant of $\rho_{v}(\tau)$, which is $4^{k-1}$. Then $\tau(\zeta)=\zeta^{4^{k-1}}$ for every $\zeta\in{\boldsymbol{\mu}}_{p^{\infty}}$.
@@
We now check Assumption 3.3. We consider the filtration given by ${\mathbb{L}}_{n}={\mathbb{Q}}(p^{n})$. Let $n\geq 1$. Fix a place of ${\mathbb{Q}}(p^{\infty})$ over $p$ corresponding to an embedding ¯¯$\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}}_{p}$. Then, by Lemma 4.7 $(i)$ and $(ii)$, and in the notation of Assumption (3.3), $e_{n}=\delta q^{n-1}=\frac{(q-1)q^{n-1}}{(q-1,k-1)}$, $i_{n}=\frac{q^{n-1}-1}{(q-1,k-1)}$ and $H_{n}\simeq\mathrm{Gal}(\mathbb{Q}_{q}(p^{n})/\mathbb{Q}_{q}(p^{n-1}))$. Thus
$\frac{e_{n}}{i_{n}+1}=\frac{(q-1)q^{n-1}}{q^{n-1}-1+(q-1,k-1)}\leq q-1.$
By Proposition 5.1, ${\overline{H_{n}}}=\mathrm{Gal}({\mathbb{L}}_{n}/{\mathbb{L}}_{n-1})$, so that ${\mathbb{L}}_{n}^{\overline{H_{n}}}={\mathbb{L}}_{n-1}$. Moreover, by Remark 5.2, $|{\overline{H_{n}}}|\leq p^{4[{\mathbb{K}}:{\mathbb{Q}}]}$.
The Assumption 3.3 therefore holds with $C_{1}=p^{2}$ and $C_{2}=p^{4[\mathbb{K}:\mathbb{Q}]}$. ∎
Proof of Theorem 1.5.
Let again ${\mathbb{F}}=\mathbb{Q}(\rho^{(p)})$. Then ${\mathbb{F}}$ is unramified at $p$, and it has Property (B) by Proposition 6.1. By Theorem 1.4, ${\mathbb{F}}(\rho_{p})$ has Property (B). But
${\mathbb{F}}(\rho_{p})=\mathbb{Q}(\rho^{(p)}\times\rho_{p})=\mathbb{Q}(\rho).$
@@ Appendix A On the linear groups of a finite product of local rings
Let ${\mathcal{A}}=\prod_{i=1}^{r}{\mathcal{A}}_{i}$ be a finite product of local commutative $\mathbb{F}_{p}$-algebras. In this appendix we bring together some results on the linear groups overs ${\mathcal{A}}$.
@@ A.1. A result on the normal closure
Let $p$ be a prime number and $k$ be a positive integer. We define
$\widehat{G}({\mathcal{A}})=\{\gamma\in\mathrm{GL}_{2}({\mathcal{A}})\ |\ \det(\gamma)\in(\mathbb{F}_{p}^{\times})^{k-1}\}$
It is the semidirect product $\mathrm{SL}_{2}({\mathcal{A}})\rtimes(\mathbb{F}_{p}^{\times})^{k-1}$, where $x\in(\mathbb{F}^{\times})^{k-1}$ acts on $\mathrm{SL}_{2}({\mathcal{A}})$ by conjugation by $\left(\begin{smallmatrix}x&0\\ 0&1\end{smallmatrix}\right)$. The aim of this subsection is to prove the following proposition.
@@
Let us assume $p\geq 5$. Let $s=\frac{p-1}{(k-1,p-1)}$ and let $H$ be a subgroup of $\mathrm{GL}_{2}(\mathbb{F}_{p})$ such that $|H|>2s$ and such that the determinant maps surjectively onto $(\mathbb{F}_{p}^{\times})^{k-1}$. Then the normal closure of $H$ in $\widehat{G}({\mathcal{A}})$ is $\widehat{G}({\mathcal{A}})$ itself.
@@
Let ${\mathcal{B}}$ be the $\mathbb{F}_{p}$-vector subspace of ${\mathcal{A}}$ generated by the squares of all the invertible elements of ${\mathcal{A}}$. Note that ${\mathcal{B}}$ is a subring with unity of ${\mathcal{A}}$ which contains $\mathbb{F}_{p}$. For every $i=1,\ldots,k$ denote by $e_{i}\in{\mathcal{A}}$ the idempotent projecting to the summand ${\mathcal{A}}_{i}$. Firstly we show that $e_{i}\in{\mathcal{B}}$ for every $i$. Indeed, choose two elements $a$, $b$ in $\mathbb{F}_{p}^{\times}$ such that $a^{2}\neq b^{2}$ (which is possible because $|\mathbb{F}_{p}^{\times}|>2$) and let $\alpha=ae_{i}+\sum_{j\not=i}e_{j}$, $\beta=be_{i}+\sum_{j\not=i}e_{j}$. Then $\alpha^{2}-\beta^{2}=(a^{2}-b^{2})e_{i}$ is a non-zero $\mathbb{F}_{p}$-multiple of $e_{i}$ belonging to ${\mathcal{B}}$.
@@
Let us assume $p\geq 5$. Let $s=\frac{p-1}{(k-1,p-1)}$ and let $H$ be a subgroup of $\widehat{G}(\mathbb{F}_{p})$ such that $|H|>2s$ and such that the determinant maps surjectively onto $(\mathbb{F}_{p}^{\times})^{k-1}$. Then the normal closure of $H$ in $\widehat{G}(\mathbb{F}_{p})$ is $\widehat{G}(\mathbb{F}_{p})$ itself.
@@
Let $N$ be the normal closure of $H$ in $\widehat{G}(\mathbb{F}_{p})$. Since $|N|>2s$, the intersection $N\cap\mathrm{SL}_{2}(\mathbb{F}_{p})$ is a non trivial normal subgroup of $\mathrm{SL}_{2}(\mathbb{F}_{p})$ of order strictly greater than $2$. Since $|\mathbb{F}|\geq 4$, the only proper normal subgroup of $SL_{2}(\mathbb{F})$ is {±$\{\pm Id\}$ (see for example [29, Theorem 8.3 and Lemma 8.6, p.539]). Thus $\mathrm{SL}_{2}(\mathbb{F}_{p})\subseteq N$. Since the projection over $(\mathbb{F}_{p}^{\times})^{k-1}$ is surjective, the result follows. ∎
We can now prove Proposition A.1.1. Let $N$ be the normal closure of $H$ in $\widehat{G}({\mathcal{A}})$. By Lemma A.1.4, $N$ contains $\widehat{G}(\mathbb{F}_{p})$, so that it contains $\mathrm{SL}_{2}({\mathcal{A}})$ by Lemma A.1.3. The surjectivity of the determinant then shows that $N=\widehat{G}({\mathcal{A}})$.
@@ A.2. Irreducibility of the adjoint representation
The argument is standard. By (A.1), any invariant subspace containing one of $A,B,C$ will also contain the others. Given some arbitrary (non-zero) matrix $M=aA+bB+cC$ of trace 0, put
@@
We can now prove Proposition A.2.5. Firstly we notice that $\mathrm{M}_{2}(\mathbb{F}_{p})^{0}\subseteq V$ by Lemma A.2.6. Since $\mathrm{M}_{2}({\mathcal{A}})^{0}=\mathrm{M}_{2}(\mathbb{F}_{p})^{0}\otimes_{\mathbb{F}_{p}}{\mathcal{A}}$, it suffices to show that $V$ is closed under the multiplication by elements in ${\mathcal{A}}$. Let $\alpha$ be an invertible element of ${\mathcal{A}}$ and let $B=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$. We have
@@
Acknowledgements
We are grateful to Andrea Conti, Samuel Le Fourn and Davide Lombardi for helpful discussions.
We thank the INdAM GNSAGA for support. We acknowledge the IEA (International Emerging Action) project PAriAIPP (Problèmes sur l’Arithmétique et l’Algèbre des Petits Points) of the CNRS for the financial support.